[figure]

4 4

4 4

Der Herr segne und behüte dich

Der

Herr

[figure]

seg _{- ne} _{und} _{be}

- hü _{- te}

[figure] [figure]

dich, [figure] [figure]

[figure]

und laß

[figure]

[figure]

leuch_{ten} [figure] [figure] ü [figure] [figure]

- ber _{dir} _{sein}

[figure: 4; An; - ge; und; sei; dir; - sicht; gnä; -  - dig; und; sei; dir; und; sei; dir; gnä; - dig; gnä; -  - dig,; der; und; sei; dir; gnä; - dig; der]
[figure: 7; Herr; er; - he; - be; ü; - ber; dir; sein; An; - ge; - sicht; und; geb; dir; Frie; geb; dir; Frie; -den; -  - den; geb; dir]
[figure: 10; Frie; geb; dir; Frie; den,; - den; dir; Frie; den,; Der; Herr; seg; -ne; und; be; - hü; - te; A; dich; p; T; ???; -  -; -  -]
[figure: 15; men,; A; ment,; T; p; Hey; -  - men,; A; -  -  - ment,; T; -  - man,; Hey; mp; Ach; -  - men,; A; -  - ment,; T; -  -  -; - man,; cresc.; -  - mens; Ach; -  - ment,; Hey; - man,; f; -  -  -  -  -; -  -  -  -  -]
[figure: 19]

men, [figure] T

dim. poco a poco rit.

mens,

[figure] Hey [figure]

- man, A [figure] Ach [figure]

- men, [figure]

- mens, [figure]

[figure]  [figure]

Hey [figure]

-  -  -  - ment,

- man, A [figure] Ach [figure]

[figure] _{}_{} T_{}_{}_{}_{}[figure] _{}_{}

Hey - - mens,

-  -  -  - men,

-  -  - ment,

Ach [figure]

A T [figure]

[figure: - men.]

- ment.

-  -  -  -  - man.
[figure: pp; -  - mens.]